Cost Reduction Proposal: IRR, NPV, and Payback Period JB Chemical currently discharges liquid waste into Calgarys municipal(a) Determine the projects net present value at a discount rate of 26 percent. (Round to the nearest whole number. Use a neg
Answer #1

Part (a)

Calculation of savings of surcharge for the year if proposed system is installed:

Surcharge per day = (200,000 litres/1000 litres * $ 4) = $ 800 per day

Cost for the year = $ 800 * 300 days = $ 240,000

Annual operating cost if the new system is installed =$ 150,000

Thus, Net savings from the installation of new system = 240,000 - 150,000

= $ 90,000

Calculation of NPV from the proposed system

Year Amount PVF @ 26% Present Value
Initial Cost 0 -360,000 1 -360,000
Net savings 1-12 90,000 3.6059 324,531
Salvage Value 12 40,000 0.0625 2,500
NPV -32969

Project's NPV is -$ 32,969.

Part (b)

IRR would be the discounting rate at which NPV = 0

Solving for IRR using scientific/financial calculator

CF0 = - 360,000

CF1-11 = 90,000

CF12 = 130,000

IRR: CPT = 23.1584%

Note: The same can also be done using trial & error method. The answer would still be the same

Part (c)

Payback period refers to the length of time required for an investment to recover its initial outlay in terms of profits or savings.

In the given case, the NPV of the proposed system is negative, i.e. the project is unable to recover it's initial outlay. Hence, the project will not have a payback period. This means that the company will not be able to recover it's initial outlay at a discounting rate of 26%.